# ST-Booster: An Iterative SpatioTemporal Perception Booster for Vision-and-Language Navigation in Continuous Environments

### v1

@@ III-D Value-Guided Waypoint Generation Module
III-E Pretraining and Fine-Tuning
Pretraining. To enhance the representational capability of the hybrid multi-modal maps, the following three pretraining tasks are introduced.
1) Masked Language Modeling (MLM). This task is commonly used to improve the model’s understanding of complex language structures [38]. To be specific, the words of the input instructions are randomly masked with a probability of 15%, prompting the model to predict the masked words based on contextual information. For MLM task, the prediction loss for generating the masked words is defined by
$\mathcal{I}_{\textrm{MLM}}=-\mathbb{E}_{(\mathbf{I},\mathbf{\Omega})\sim\mathcal{D}}\log\mathcal{P}_{\bm{\theta}}(\mathbf{I}_{\text{mask}}|\mathbf{I}_{\setminus\text{mask}},\mathbf{G}_{t},\mathbf{M}_{t}),$ (8)
where $\bm{\theta}$ is the policy parameters, $\mathbf{I}_{\text{mask}}$ and $\mathbf{I}_{\setminus\text{mask}}$ represent the masked and the unmasked portion of the instruction embeddings, respectively, and $\mathbf{\Omega}$ denotes the expert trajectory.
2) Hybrid Single Action Prediction (HSAP). This task aims to enhance the model’s environmental perception and instruction comprehension by improving decision-making accuracy [18]. Specifically, we utilize $\mathbf{G}_{t}$ and $\mathbf{M}_{t}$ to predict the long-term waypoint-selected action $\mathbf{a}^{G}_{t}$ and short-term waypoint-selected action $\mathbf{a}^{M}_{t}$, respectively. Then, we fuse them to get a more robust decision. Specifically, if $\mathbf{a}^{G}_{t}$ is among the current candidate waypoint set $\mathcal{V}^{\text{o}}_{t,\text{cur}}$, we perform a weighted fusion of $\mathbf{a}^{G}_{t}$ and $\mathbf{a}^{M}_{t}$ with the weight $\gamma_{t}$; otherwise, we directly adopt $\mathbf{a}^{G}_{t}$, i.e.,
$\mathbf{a}_{t}^{G}=\textrm{FFN}(\tilde{\mathbf{G}}_{t}),\quad\mathbf{a}_{t}^{M}=\textrm{FFN}(\tilde{\mathbf{M}}_{t}),$ (9)
$\mathbf{a}_{t}=\left\{\begin{aligned} &\mathbf{\gamma}_{t}\mathbf{a}_{t}^{G}+(1-\mathbf{\gamma}_{t})\mathbf{a}_{t}^{M},~{}\textrm{if}~{}\mathbf{a}_{t}^{G}\in\mathcal{V}^{\text{o}}_{t,\text{cur}},\\ &\mathbf{a}_{t}^{G},~{}\text{otherwise}.\end{aligned}\right.$ (10)
@@ III-E Pretraining and Fine-Tuning
$\mathcal{L}_{\textrm{HSAP}}=-\mathbb{E}_{(\mathbf{L},\mathbf{\Omega},{\mathbf{a}_{t}^{*}})\sim\mathcal{D}}\log\mathcal{P}_{\bm{\theta}}(\mathbf{a}_{t}^{*}|\mathbf{I},\mathbf{G}_{t},\mathbf{M}_{t}).$ (11)
@@
$\mathcal{L}_{\textrm{GAHP}}=\mathbb{E}_{\mathbf{H}_{t}^{*}\sim\mathcal{D}}(\textrm{MSE}(\mathbf{H}_{t}^{*},\mathbf{H}_{t})|\mathbf{\tilde{M}}_{t}).$ (12)
TABLE I: Performance comparison of our proposed ST-Booster against state-of-the-art methods evaluated on the VLN-CE dataset. The bolded numbers represent the optimal performance. The underlined numbers indicate the second-best performance)
[…]
Seq2Seq [3] 9.26 7.12 46 37 35 8.64 7.37 40 32 30 8.85 7.91 36 28 25
CM2 [23] 12.05 6.10 51 43 35 11.54 7.02 42 34 28 13.90 7.70 39 31 24
MGMap [24] 10.12 5.65 52 47 43 10.00 6.28 48 39 34 12.30 7.11 45 35 28
CWP-RecBERT [21] 12.50 5.02 59 50 44 12.23 5.74 53 44 39 13.31 5.89 51 42 36
Sim2Sim [22] 11.18 4.67 61 52 44 10.69 6.07 52 43 36 11.43 6.17 52 44 37
ETPNav [38] 11.78 3.95 72 66 59 11.99 4.71 65 57 49 12.87 5.12 63 55 48
Ego2-Map [41] 4.94 52 46 13.05 5.54 56 47 41
GridMM [7] 12.69 4.21 69 59 51 13.36 5.11 61 49 41 13.31 5.64 56 46 39
BEVBert [18] 4.57 67 59 50 4.70 67 59 50
[…]
Fine-Tuning. To prevent the model from overfitting to the offline expert data, we fine-tune the pre-trained model in the continuous simulation environments with the HSAP task. During this fine-tuning process, the topological and grid maps are updated in real-time based on the agent’s observations. Additionally, we progressively transition from teacher-forcing to student-forcing during action reasoning phase [21] to balance learning stability and exploration. Specifically, in teacher-forcing strategy, the agent navigates with ground-truth actions, enabling the model to quickly learn reasonable paths toward the target, while in student-forcing strategy, the agent explores the unknown state-action space by navigating based on sampled results from the model’s output This hybrid training approach effectively balances exploration and exploitation, allowing the model to learn expert behaviors effectively.
@@ IV Experiments
This section initially presents the experimental setup, detailing the dataset, training implementation, and evaluation metrics. Subsequently, systematic comparisons are conducted between ST-Booster and state-of-the-art baselines in VLN-CE. Through extensive experiments and analysis, we demonstrate the functional contributions of each core module in ST-Booster.
@@ IV-A Experimental Setup
Datasets. The Matterport3D (MP3D) environment [3] with 90 scenes and 10,800 RGB-D images is commonly used for discrete VLN tasks. To enable continuous navigation, the VLN-CE benchmark reconstructs these scenes in 3D within the Habitat simulator [43], adapting the Room-to-Room (R2R) dataset [3] for continuous environments. The dataset is split into train, val seen, val unseen, and test subsets. The val seen split involves new paths and instructions within familiar training environments, while val unseen and test contain paths in novel, unseen environments. To evaluate the proposed framework’s generalization capabilities, models are typically assessed on unseen environments.
[…]
Evaluation Metrics. To evaluate the performance of the proposed model, we utilized several standard metrics, including (i) Trajectory Length (TL), measuring the average path length; (ii) Navigation Error (NE), representing the average distance between the agent’s final position and the target; (iii) Oracle Success Rate (OSR), the rate at which the agent reaches a position within 3m of the target; (iv) Success Rate (SR), the rate at which the agent stops within 3m of the target; and (v) Success rate weighted by Path Length (SPL), which balances success rate with path efficiency. These metrics provide a comprehensive assessment of the model’s navigation performance on both seen and unseen splits.
Figure 4: Successful navigation examples with short and long instructions in unseen environments.The figure presents four representative examples. The upper two groups demonstrate navigation scenarios corresponding to concise textual instructions (displayed above the navigation maps), featuring the agent’s real-time panoramic observations and top-down views during navigation. The lower two groups illustrate navigation scenarios guided by extended textual instructions.
IV-B Comparison with the State-of-the-Art Methods
We conducted comprehensive comparisons with several state-of-the-art approaches listed in Table I. For instance, Sim2Sim [22] and CWP-RecBERT [21] employ the end-to-end training framework to predict actions through joint modeling of vision features and historical latent variables. ETPNav [38] utilizes topological map encoding to capture historical navigation patterns and derive optimal planning trajectories. GridMM [7] implements a grid-based representation mechanism to model fine-grained environmental details for precise navigation. The method with the most comparable performance to ours is BEVBert [18], which is a hybrid map-based method. BEVBert employs separate training procedures for different map representations and then fuses the outputs from the two maps during the navigation planning phase. In contrast, ST-Booster integrates a hybrid map feature transfer and fusion mechanism. The enhanced perception is utilized to optimize both the waypoint generation and the navigation planning processes.
As illustrated in Table I, for the val unseen split, ST-Booster improves the highest SR (BEVBert) by 2%, indicating its superior generalization to unseen environments. In the test-unseen split, ST-Booster achieves higher SR (59%) and SPL (50%) compared to competing approaches. Furthermore, as demonstrated in Figure 4, several visualization examples with short and long instructions illustrate the navigation processes in unseen environments using ST-Booster. These examples highlight that ST-Booster, by enhancing multi-modal hybrid map perception, effectively captures spatiotemporal relationships, and improves navigation task success rates.
@@ IV-B Comparison with the State-of-the-Art Methods
[…]
@@
[…]
@@
Figure 6: Generated GAHs of BEVBert and ST-Booster on val unseen split of VLN-CE. (a) Generated GAHs of BEVBert and ST-Booster on val unseen split of VLN-CE. (b) Ground-truth and predicted GAHs examples along the paths on VLN-CE val unseen split. The experimental results demonstrate that the GAHs generated by ST-Booster effectively and accurately highlight the key regions associated with landmarks in the instructions.
Figure 7: Robustness performance analysis after introducing disturbances using different VLN-CE agents ST-Booster (ST) and BEVBert (BEV). (a) SR under different waypoint disturbance ratio. ‘BEV+GAH’ means the BEVBert agent with VGWG that generats candidate waypoints from GAHs. (b) SPL under different waypoint disturbance ratio. (c) SR under different perception disturbance ratio. ‘VLN-CE Global’ means the VLN-CE (ST/BEV) agent with the global perception disturbance. ‘VLN-CE Local’ means the VLN-CE agent with the local perception disturbance. (d) SPL Under Different Perception Disturbance Ratio.
@@ IV-C Ablation Study
This ablation study aims to validate the synergistic optimization effects of the feature fusion capability of the MGAF module and the instruction-guided waypoint prediction mechanism of the VGWG module. As shown in Table II, we consider three different map experts to predict actions with or without the key modules proposed in ST-Booster, including topological, grid, and hybrid map experts. Specifically, when activating MGAF without VGWG, it improves the success rate (SR) of the topological map expert by 1.9% and the grid map expert by 3.4%. With VGWG enabled, MGAF further boosts these improvements to 2.8% and 4.6% respectively, while achieving a 2.0% SR enhancement for the hybrid map expert. These improvements confirm MGAF’s effectiveness in facilitating complementary fusion of dual-map perceptions and strengthening instruction-feature alignment, thereby benefiting both individual and hybrid experts. Furthermore, the VGWG module demonstrates its distinct value through instruction-aware waypoint distribution prediction, which focuses attention on navigation-critical regions, contributing an additional 1.2% SR gain to the hybrid expert. The combined application of both modules enables the hybrid expert to achieve optimal performance with 61.0% SR, illustrating the effectiveness of the modules within ST-Booster.
@@ IV-D Performance Analysis
Consistency Against Task Complexity. We evaluate the generalization performance of ST-Booster in complex tasks within unseen environments, conducting a statistical analysis of the proposed framework’s performance. Task complexity is quantified based on the instruction’s complexity across four dimensions. For instance, we count the number of terms representing spatial elements and landmarks in each instruction, where a higher count indicates increased spatial complexity of the navigation task. Similarly, we analyze the frequency of numeric terms and the total vocabulary count in each instruction, reflecting the task’s cognitive difficulty in handling long-term dependencies. Experimental results in Figure 5 demonstrate that the performance of ST-Booster is consistently superior to the selected SOTA methods across all complexity dimensions. This confirms that ST-Booster effectively enhances complex spatiotemporal perception, and improves navigation performance.
Visualization of Predicted GAHs. To visually validate ST-Booster’s enhanced environmental perception, we conduct a comparative analysis of GAHs generated from both our augmented perceptual features and baseline image features. As shown in Figure 6 (a), the GAHs generated using ST-Booster effectively highlight the key regions associated with landmarks in the instructions. This indicates that leveraging the ST-Booster-enhanced environmental perception features enables improved environmental understanding and instruction alignment, capturing contextual cues from the local environment more effectively. In Figure 6 (b), we illustrate some comparisons between ground-truth and predicted GAHs during evaluation in unseen environments. In the presented examples, the ground-truth GAHs accurately reflect instruction-relevant regions in the current surroundings, while the predicted GAHs also provide similar cues to some extent. However, since GAHs represent coarse-grained information, we discourage assigning it excessive weight during utilization, as this may introduce additional errors and disrupt the decision-making process. Notably, we do not recommend using GAHs during the fine-tuning stage, as it may constrain the model’s exploration of the environment, limiting the learning performance.
Robustness Against Disturbance. To simulate environmental interference commonly encountered in real-world scenarios, we implement several distinct types of disturbances, followed by comprehensive performance analysis. (i) Waypoint Generation Disturbance. The first experiment emulates visual perception degradation by probabilistically substituting waypoint predictions with random sampling. As shown in Figure 7 (a) and (b), our framework demonstrates waypoint disturbance tolerance, maintaining 59% SR and 47.7% SPL with waypoint disturbance ratio 9%, which is 1.5% and 1.1% higher than BEVBert baseline. This resilience originates from GAHs that compensate visual prediction errors through instruction-anchored spatial priors. (ii) Global Perception Disturbance. The second experiment evaluates hybrid perception robustness against global perception disturbance. We progressively discard historical nodes to simulate privacy-constrained memory management. As revealed in Figure 7 (c) and (d), our dual-map architecture maintains 58.75% SR and 48.68% SPL performance with global disturbance ratio 90%, outperforming the baseline by 1.28% and 1.37%, respectively. (iii) Local Perception Disturbance. In addition, we also valuate hybrid perception robustness against local perception disturbance by applying Gaussian blur kernels and random visual masking to mimic motion artifacts and occlusions. With local disturbance ratio 90% in Figure 7 (c) and (d), our method outperforms better than the baseline with 1.41% SR and 2.94% SPL. This robustness stems from cross-domain complementarity, where topological graphs preserve structural consistency through geometric detail integration from grids, while grid maps retain local accuracy via global context propagation from nodes.

### v2

@@ 20 $\mathbf{P}_{t+1}=\text{Add}(\mathbf{P}_{t},\mathbf{O}_{t+1})$
1) Masked Language Modeling (MLM). This task is commonly used to improve the model’s understanding of complex language structures [an2024etpnav]. To be specific, the words of the input instructions are randomly masked with a probability of 15%, prompting the model to predict the masked words based on contextual information. For the MLM task, the prediction loss for generating the masked words is defined by
$\mathcal{I}_{\textrm{MLM}}=-\mathbb{E}_{(\mathbf{I},\mathbf{\Omega})\sim\mathcal{D}}\log\mathcal{P}_{\bm{\theta}}(\mathbf{I}_{\text{mask}}|\mathbf{I}_{\setminus\text{mask}},\tilde{\mathbf{G}}^{\text{f}}_{t},\tilde{\mathbf{M}}^{\text{f}}_{t}),$ (8)
where $\bm{\theta}$ denotes the policy parameters, $\mathbf{I}_{\text{mask}}$ and $\mathbf{I}_{\setminus\text{mask}}$ represent the masked and the unmasked portion of the instruction embeddings, respectively, and $\mathbf{\Omega}$ denotes the expert trajectory.
2) Hybrid Single Action Prediction (HSAP). To enhance the environmental perception and instruction comprehension by improving decision-making accuracy [an2023bevbert], $\tilde{\mathbf{G}}^{\text{f}}_{t}$ and $\tilde{\mathbf{M}}^{\text{f}}_{t}$ are utilized to predict the long-term action $\mathbf{a}^{\text{G}}_{t}$ and short-term action $\mathbf{a}^{\text{M}}_{t}$, respectively. Then, we fuse them to get a robust decision. Specifically, if $\mathbf{a}^{\text{G}}_{t}$ is among the current candidate waypoint set $\mathcal{V}^{\text{o}}_{t,\text{cur}}$, we perform a weighted fusion of $\mathbf{a}^{\text{G}}_{t}$ and $\mathbf{a}^{\text{M}}_{t}$ with the weight $\gamma_{t}$; otherwise, we directly adopt $\mathbf{a}^{\text{G}}_{t}$, i.e.,
$\mathbf{a}_{t}^{\text{G}}=\textrm{FFN}(\tilde{\mathbf{G}}^{\text{f}}_{t}),\quad\mathbf{a}_{t}^{\text{M}}=\textrm{FFN}(\tilde{\mathbf{M}}^{\text{f}}_{t}),$ (9)
$\mathbf{a}_{t}=\left\{\begin{aligned} &\mathbf{\gamma}_{t}\mathbf{a}_{t}^{\text{G}}+(1-\mathbf{\gamma}_{t})\mathbf{a}_{t}^{\text{M}},~\textrm{if}~\mathbf{a}_{t}^{\text{G}}\in\mathcal{V}^{\text{o}}_{t,\text{cur}},\\ &\mathbf{a}_{t}^{\text{G}},~\text{otherwise}.\end{aligned}\right$. (10)
@@
$\mathcal{L}_{\textrm{HSAP}}=-\mathbb{E}_{(\mathbf{L},\mathbf{\Omega},{\mathbf{a}_{t}^{*}})\sim\mathcal{D}}\log\mathcal{P}_{\bm{\theta}}(\mathbf{a}_{t}^{*}|\mathbf{I},\tilde{\mathbf{G}}^{\text{f}}_{t},\tilde{\mathbf{M}}^{\text{f}}_{t}).$ (11)
@@
Figure 3: Successful navigation examples under short and long instructions in unseen environments, along with visualizations of predicted and ground-truth GAHs along successful paths. The figure presents four representative cases. The top example illustrates navigation guided by a concise textual instruction (displayed above the navigation map), including the agent’s real-time panoramic observations and top-down views. The middle two examples depict navigation scenarios guided by longer instructions, while the bottom example shows the visualization of predicted and ground-truth GAHs along a successful trajectory.
$\mathcal{L}_{\textrm{GAHP}}=\mathbb{E}_{\mathbf{H}_{t}^{*}\sim\mathcal{D}}(\textrm{MSE}(\mathbf{H}_{t}^{*},\mathbf{H}_{t})|\tilde{\mathbf{M}}^{\text{f}}_{t}).$ (12)
Fine-Tuning. To prevent the model from overfitting to the offline expert data, we fine-tune the pre-trained model in the continuous simulation environments with the HSAP task. During this fine-tuning process, the topological and grid maps are updated in real-time based on the agent’s observations. Additionally, we progressively transition from teacher-forcing to student-forcing during action reasoning phase [hong2022bridging] to balance learning stability and exploration. Specifically, in the teacher-forcing strategy, the agent navigates with ground-truth actions, enabling the model to quickly learn reasonable paths toward the target, while in the student-forcing strategy, the agent explores the unknown state-action space by navigating based on sampled results from the model’s output This hybrid training approach effectively balances exploration and exploitation, allowing the model to learn expert behaviors effectively.
@@ IV Experiments
This section presents the experimental setup, including the dataset, training details, and evaluation metrics, followed by comparisons between ST-Booster and SoTA baselines in VLN-CE. Extensive experiments further validate the contribution of each core module.
@@ IV-A Experimental Setup
Datasets. The Matterport3D (MP3D) environment [anderson2018vision] with 90 scenes and 10,800 RGB-D images is commonly used for discrete VLN tasks. To enable continuous navigation, the VLN-CE benchmark reconstructs these scenes in 3D within the Habitat simulator [2019habitat], adapting the Room-to-Room (R2R) dataset [anderson2018vision] for continuous environments. The dataset is split into train, val seen, val unseen, and test subsets. The val seen split involves new paths and instructions within familiar training environments, while val unseen and test contain paths in novel, unseen environments. To evaluate the proposed framework’s generalization capabilities, models are typically assessed on unseen environments.
TABLE I: Performance comparison of our proposed ST-Booster against state-of-the-art methods evaluated on the VLN-CE dataset. (P-GAH / FT-GAH: pretrained and fine-tuned GAH modules; EV+: evaluation using only GAH predictions; ST-FT / ST-EV: fine-tuning and evaluation stages of ST-Booster)
Method Val Seen Val Unseen Test Unseen
TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$) TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$) TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$)
Seq2Seq [anderson2018vision] 9.26 7.12 46 37 35 8.64 7.37 40 32 30 8.85 7.91 36 28 25
CM2 [georgakis2022cross] 12.05 6.10 51 43 35 11.54 7.02 42 34 28 13.90 7.70 39 31 24
MGMap [chen2022weakly] 10.12 5.65 52 47 43 10.00 6.28 48 39 34 12.30 7.11 45 35 28
CWP-RecBERT [hong2022bridging] 12.50 5.02 59 50 44 12.23 5.74 53 44 39 13.31 5.89 51 42 36
Sim2Sim [krantz2022sim] 11.18 4.67 61 52 44 10.69 6.07 52 43 36 11.43 6.17 52 44 37
ETPNav [an2024etpnav] 11.78 3.95 72 66 59 11.99 4.71 65 57 49 12.87 5.12 63 55 48
Ego2-Map [hong2023learning] 4.94 52 46 13.05 5.54 56 47 41
GridMM [wang2023gridmm] 12.69 4.21 69 59 51 13.36 5.11 61 49 41 13.31 5.64 56 46 39
BEVBert [an2023bevbert] 4.57 67 59 50 4.70 67 59 50
Scale-VLN [wang2023scale] 4.80 55 51 5.11 55 50
ST-Booster (Ours) 13.29 3.71 77 69 59 13.14 4.77 68 61 50 13.42 4.83 66 59 50
@@
Expert MGAF VGWG TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$)
Topological Map 13.87 5.24 61.6 53.6 43.1
@@
13.14 4.77 67.6 61.0 50.2
Implementation Details. The visual feature extraction of ST-Booster is based on ViT-B/16-CLIP [radford2021learning], which generates representations with a dimension of 768. For grid-based spatial representations, the extracted features are downsampled to match the encoder patch size $P\times P=$$\times14$. When the spatial dimensions are set to $U=V=11$ with a cell resolution of 1 m, the grid map covers an area of $11\,\text{m}\times 11\,\text{m}$, where each cell corresponds to 1 $\text{m}^{2}$. With an upsampling factor of $m=n=5$, the GAH map expands to $55\,\text{m}\times 55\,\text{m}$, and each cell has a side length of 0.2 m. During pretraining, our model is initialized with pre-trained LXMERT weights [tan2019lxmert] and undergoes 200,000 iterations of offline training using a batch size of 64. The model with the highest pretraining performance is subsequently fine-tuned through an additional 20,000 iterations of online training with a batch size of 16. Finally, the fine-tuned model is evaluated on unseen environments, and the highest-performing variant is selected for further analysis.
Figure 4: Visualization of GAH generator predictions. (a) Generated GAHs of BEVBert and ST-Booster on val unseen split of VLN-CE. (b) The influence of GAH on candidate waypoint distributions under correct and erroneous predictions. (c) Successful and failed examples generated by GAH.
Evaluation Metrics. To evaluate the performance of the proposed model, we utilized several standard metrics, including (i) Trajectory Length (TL), measuring the average path length; (ii) Navigation Error (NE), representing the average distance between the agent’s final position and the target; (iii) Oracle Success Rate (OSR), the rate at which the agent reaches a position within $3\,\text{m}$ of the target; (iv) Success Rate (SR), the rate at which the agent stops within $3\,\text{m}$ of the target; and (v) Success rate weighted by Path Length (SPL), which balances success rate with path efficiency. These metrics provide a comprehensive assessment of the model’s navigation performance on both seen and unseen splits.
IV-B Comparison With the State-of-the-Art Methods
We compare ST-Booster with several state-of-the-art methods listed in Table I. Sim2Sim [krantz2022sim] and CWP-RecBERT [hong2022bridging] adopt end-to-end frameworks that jointly model visual features and historical states, while ETPNav [an2024etpnav] encodes topological maps to capture navigation patterns, and GridMM [wang2023gridmm] applies grid-based representations for fine-grained spatial reasoning. BEVBert [an2023bevbert], the most comparable hybrid-map approach, trains global and local maps separately and fuses them at the decision stage. In contrast, ST-Booster performs feature-level hybrid map fusion, jointly optimizing waypoint generation and navigation reasoning through enhanced spatiotemporal perception. As illustrated in Table I, ST-Booster surpasses BEVBert by 2% in SR on the val-unseen split and achieves the best SR (59%) and SPL (50%) on the test-unseen split. Visualization results in Fig. 3 further demonstrate that by improving multi-modal hybrid map perception, ST-Booster more effectively captures spatiotemporal relationships and achieves higher navigation success in unseen environments.
TABLE III: Performance Comparison of GAH Models at Different Training Stages on the Val-Unseen Split. (P-GAH and FT-GAH denote the GAH prediction modules obtained after pretraining and fine-tuning, respectively; EV+ indicates decision-making using only GAH predictions during evaluation; ST-FT and ST-EV represent the fine-tuning and evaluation stages of ST-Booster, respectively)
Method with GAH TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$)
EV+P-GAH 20.58 8.51 27.35 19.74 10.12
EV+FT-GAH 19.02 8.15 28.87 21.15 11.06
ST-FT+P-GAH 16.51 4.79 68.03 58.08 45.03
ST-FT+FT-GAH 15.18 4.74 65.85 58.78 47.67
ST-EV+P-GAH 13.14 4.77 67.50 60.95 50.16
ST-EV+FT-GAH 13.02 4.72 67.54 61.12 50.55
@@ IV-B Comparison With the State-of-the-Art Methods
Figure 6: Robustness analysis of different VLN-CE agents, ST-Booster (ST) and BEVBert (BEV), under various disturbances. (a) SR and path efficiency SPL under different waypoint disturbance ratios. ‘BEV+GAH’ denotes BEVBert with the VGWG module generating candidate waypoints from GAHs. (b) SR and SPL under different perception disturbance levels, where ‘VLN-CE Global’ and ‘VLN-CE Local’ indicate agents affected by global and local perception disturbances, respectively. (c) Decision consistency results, where LSR and GSR denote the SR of local and global experts, and GEL indicates the agreement probability between their decisions.
@@ IV-C Ablation Study
This ablation study aims to validate the synergistic optimization effects of the feature fusion capability of the MGAF module and the instruction-guided waypoint prediction mechanism of the VGWG module. As illustrated in Table II, we consider three different map experts to predict actions with or without the key modules proposed in ST-Booster, including topological, grid, and hybrid map experts. Specifically, when activating MGAF without VGWG, it improves the success rate (SR) of the topological map expert by 1.9% and the grid map expert by 3.4%. With VGWG enabled, MGAF further boosts these improvements to 2.8% and 4.6% respectively, while achieving a 2.0% SR enhancement for the hybrid map expert. These improvements confirm MGAF’s effectiveness in facilitating complementary fusion of dual-map perceptions and strengthening instruction-feature alignment, thereby benefiting both individual and hybrid experts. Furthermore, the VGWG module demonstrates its distinct value through instruction-aware waypoint distribution prediction, which focuses attention on navigation-critical regions, contributing an additional 1.2% SR gain to the hybrid expert. The combined application of both modules enables the hybrid expert to achieve optimal performance with 61.0% SR, illustrating the effectiveness of the modules within ST-Booster.
@@ IV-D Performance Analysis
GAH Performance. To comprehensively assess the role of the GAH, we analyze GAH via visualization, waypoint influence, decision ability, and train–test usage.
GAH Visualization Analysis. Compared with GAHs from baseline image features, BEVBert lacks global cues and produces checkerboard heatmaps (Fig. 4(a)). With enhanced representation, ST-Booster yields smoother GAHs that highlight instruction-relevant landmarks. GAH usually steers candidates toward goal regions, making the fused distribution closer to the ground truth (illustrated in the top two rows of Fig. 4(b) and Fig. 4(c)), but its coarse cues add noise in some cases (illustrated in the bottom two rows of Fig. 4(b) and Fig. 4(c)).
GAH Decision Capability Analysis. We quantitatively evaluate GAH’s decision capability. As illustrated in Table III, weighting the pretrained GAH with the candidate waypoint distribution and selecting the highest-value waypoint yields a SR of 19.74%. After fine-tuning VGWG with the mean squared error between the GAH and the ground-truth heatmap, SR rises to 21.15%. Although GAH effectively highlights instruction-relevant regions, its SR remains lower than when combined with the dual-map expert. The gap mainly stems from prediction errors and the lack of backtracking. Therefore, GAH used as an auxiliary prior to complement spatial information improves the robustness of the dual-map experts.
GAH Performance Improvement Analysis. We investigate when to apply GAH. As illustrated in Table III, using GAH during fine-tuning reduces SR to 58.78% due to over-concentrated candidates and reduced exploration. In contrast, applying GAH only during evaluation improves SR to 61.12%, as the fine-tuned dual-map experts tolerate imperfect GAHs and focus on critical regions using prior knowledge. Overall, GAH may constrain training exploration but improves inference, so using it as an auxiliary module at evaluation works best.
@@
Consistency Against Task Complexity. We evaluate the generalization performance of ST-Booster in complex tasks within unseen environments, conducting a statistical analysis of the proposed framework’s performance. Task complexity is quantified based on the instruction’s complexity across four dimensions. For instance, we count the number of terms representing spatial elements and landmarks in each instruction, where a higher count indicates increased spatial complexity of the navigation task. Similarly, we analyze the frequency of numeric terms and the total vocabulary count in each instruction, reflecting the task’s cognitive difficulty in handling long-term dependencies. Experimental results in Fig. 5 demonstrate that the performance of ST-Booster is consistently superior to the selected SOTA methods across all complexity dimensions. This confirms that ST-Booster effectively enhances complex spatiotemporal perception and improves navigation performance.
Robustness from Spatiotemporal Complementarity. We evaluate the model’s robustness by testing the complementary effects of spatiotemporal perception under three controlled disturbances: (1) field-of-view loss, simulated by randomly selecting candidate waypoints to disrupt the uniformity of the original waypoint distribution; (2) local noise, introduced by applying motion blur and partial masking to local observations to mimic real camera disturbances; and (3) long-term memory decay, simulated by randomly removing previously visited waypoints to emulate memory limitations. As illustrated in Fig. 6, our model consistently outperforms BEVBert in SR and SPL across all conditions and shows smaller performance drops under unreliable perception. This robustness arises from the spatiotemporal complementarity between the dual-map experts. When short-term local observations become noisy, the topological expert offers long-term semantic stability. Conversely, when global topology weakens, the grid expert preserves spatial accuracy through real-time sensing. Their interaction at the feature level allows information to flow and adjust across time and space, maintaining stable navigation under various disturbances.
Consistency from Spatiotemporal Alignment. We evaluate decision consistency achieved through spatiotemporal alignment between the two experts. In BEVBert, the grid and topological experts act independently and are fused only at the decision stage, leading to temporal and semantic misalignment. As illustrated in Fig. 6(c), our model enables bidirectional spatiotemporal feature fusion, aligning local observations with global semantics across time and space. This alignment improves information integration, combining short-term perception with long-term memory for more coherent decisions. In unseen environments, it increases SR of the grid and topological experts to 49% and 57%, respectively, and raises decision agreement to 58%, a 6% gain over BEVBert, confirming more consistent and reliable navigation.
TABLE IV: The effect of grid map size on navigation performance
Scale Size Cell Resolution TL NE ($\downarrow$) OSR ($\uparrow$) SR ($\uparrow$) SPL ($\uparrow$)
$\times 7$ 0.5 m 14.02 5.10 63.9 55.6 45.6
1 m 12.73 4.80 63.9 57.2 47.9
$\times11$ 0.5 m 13.05 4.99 65.7 57.9 49.2
1 m 13.14 4.77 67.6 61.0 50.2
$\times15$ 0.5 m 12.70 4.81 66.2 58.7 49.9
1 m 13.2 4.78 66.6 59.8 48.8